OpenStudy (anonymous):

I think I am setting up the equation wrong..? Question is: A park in a subdivision is triangular-shaped. Two adjacent sides of the park are 533 feet and 525 feet. The angle between the sides is 53 degrees. Find the area of the park to the nearest square foot. I thought this was what I was suppose to do. 1/2 * 533 * 525 * sin (53) But I keep getting a HUGE answer. So I don't think that's it. And I think I am suppose to divide after that but I am not for sure what I should be dividing by. Please help!

OpenStudy (anonymous):

@geerky42 Hey could you help again? :)

geerky42 (geerky42):

Hmm, I don't see anything wrong with it.

geerky42 (geerky42):

I think you did it correctly.

geerky42 (geerky42):

Something answer may be hugely different than you except. :)

OpenStudy (anonymous):

Do I need to divide by anything?

geerky42 (geerky42):

no, just \(\dfrac{1}{2}\cdot533 \cdot 525 \cdot\sin (53^\text{o})\)

OpenStudy (anonymous):

Maybe I am entering it into my calculator wrong? This is the answer I am getting -5883002.6088453643...

geerky42 (geerky42):

Make sure your calculator is in degree mode.

OpenStudy (anonymous):

Yeah, it is.

geerky42 (geerky42):

Hmm, calculate \(\sin (53º)\)

geerky42 (geerky42):

what do you get?

OpenStudy (anonymous):

0.798635510047...

geerky42 (geerky42):

hmm same here. now multiply it by 0.5*533*525

OpenStudy (anonymous):

Now I got 111739.0907994918..

geerky42 (geerky42):

Now, this is correct answer (what I get.)

OpenStudy (anonymous):

I must have put it in the calculator wrong somehow. I only need 11739.09, right?

geerky42 (geerky42):

111739.09, yeah.

OpenStudy (anonymous):

Okay. Thank you so much again!! You have been SO much help!! :D

geerky42 (geerky42):

lol, glad I helped. :)